Tell whether the graphs of the pair of equations are parallel, perpendicular or neither.
5 x + 9 y = 9
9 x - 5 y = 4
I am struggling with these. Thank you for your help.
Emily

Answer by scott8148(6628)

solving for y puts the equations into slope-intercept form
___ if the slopes are the same, the graphs are parallel
___ if the slopes are negative reciprocals, the graphs are perpendicular

5 x + 9 y = 9 ___ subtracting 5x ___ 9y = -5x + 9
___ dividing by 9 ___ y = -5/9 x + 1


9 x - 5 y = 4 ___ subtracting 9x ___ -5y = -9x + 4
___ dividing by -5 ___ y = 9/5 x - 4/5

the slopes are negative reciprocals so the graphs are perpendicular